# Integrated Sensing and Communications for Pinching-Antenna Systems (PASS)

### v2

@@ II-B Problem Formulation
In this paper, we aim to maximize the illumination power $P(\theta_{\text{s}},r_{\text{s}})$ by jointly designing the pinching beamformer, under the transmit power budget and communication QoS requirement, which is given by
$(\text{P}1)\quad$ $\max\limits_{\mathbf{h}_{\text{s}},\mathbf{h}_{\text{c}},\mathbf{w}}P(\theta_{\text{s}},r_{\text{s}})$ (8a)
s.t. $|\mathbf{w}_{[n]}|^{2}=\frac{P_{\text{max}}}{N},$ (8b)
$R\geq R_{\text{QoS}},$ (8c)
$\mathbf{h}_{[n],\text{s}}=\mathcal{H}_{\text{s}}(n),$ (8d)
$\mathbf{h}_{[n],\text{c}}=\mathcal{H}_{\text{c}}(n),$ (8e)
$|x_{n}^{\text{p}}-x_{m}^{\text{p}}|\geq\Delta x,\ n\neq m,$ (8f)
$x_{n}^{\text{p}}\in[x_{n}^{\text{p,min}},x_{n}^{\text{p,max}}],\ n\in[1,N],$ (8g)
where $\mathcal{H}_{\text{s}}(n)=\frac{\eta^{\frac{1}{2}}e^{\jmath\frac{2\pi}{\lambda}\sqrt{r_{\text{s}}^{2}-2r_{\text{s}}\cos\phi_{\text{s}}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+d^{2}}}}{\sqrt{r_{\text{s}}^{2}-2r_{\text{s}}\cos\phi_{\text{s}}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+d^{2}}}$, $\mathcal{H}_{\text{c}}(n)=\frac{\eta^{\frac{1}{2}}e^{\jmath\frac{2\pi}{\lambda}\sqrt{x_{1}^{2}-2x_{1}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+y_{1}^{2}+d^{2}}}}{\sqrt{x_{1}^{2}-2x_{1}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+y_{1}^{2}+d^{2}}}$, $P_{\text{max}}$ denotes the maximal transmit power at the BS, and $R_{\text{QoS}}$ denotes the minimum tolerated achievable rate of the communication user. (8b) denotes the maximal transmit power constraint; (8c) represents the QoS requirement; (8f) and (8g) limits the activation range of each pinching antenna. Notably, the problem (P1) is challenging to solve due to the quadratic objective function and the coupling between the $\{\mathbf{h}_{\text{s}},\mathbf{h}_{\text{c}}\}$ and $\{\mathbf{w}\}$.
@@ III Pinching Beamforming Optimization
In this section, we focus on the pinching beamforming design by jointly optimizing the radiation pattern and activation position of pinching antennas. Since all the channels $\{\mathbf{h}_{\text{s}},\mathbf{h}_{\text{c}}\}$ are relative to $\mathbf{x}^{\text{p}}$, which denotes the x-axis coordinate of the transmitting pinching antennas. Thus, the problem (P1) can be rewritten as
$(\text{P}2)\ \max\limits_{\mathbf{x}^{\text{p}},\mathbf{w}}\quad$ $\mathbf{h}_{\text{s}}^{H}\mathbf{w}\mathbf{w}^{H}\mathbf{h}_{\text{s}}$ (9a)
s.t. $|\mathbf{h}_{\text{c}}^{H}\mathbf{w}|^{2}\geq\gamma_{\text{QoS}}\sigma^{2},$ (9b)
@@
where $\gamma_{\text{QoS}}=2^{R_{\text{QoS}}}-1$. To address the quadratic objective and constraints, we apply the SDR technique to rewrite the problem (P2) as follows.
$(\text{P}3)\ \max\limits_{\mathbf{x}^{\text{p}},\mathbf{W}}$ $\text{Tr}(\mathbf{H}_{\text{s}}\mathbf{W})$ (10a)
s.t. $\mathbf{W}_{[n,n]}=\frac{P_{\text{max}}}{N},$ (10b)
$\text{Tr}(\mathbf{H}_{\text{c}}\mathbf{W})\geq\gamma_{\text{QoS}}\sigma^{2},$ (10c)
$\mathbf{W}\succeq\mathbf{0},$ (10d)
$\text{rank}(\mathbf{W})=1,$ (10e)
@@
where $\mathbf{W}=\mathbf{w}\mathbf{w}^{H}$, $\mathbf{H}_{\text{s}}=\mathbf{h}_{\text{s}}\mathbf{h}_{\text{s}}^{H}$, and $\mathbf{H}_{\text{c}}=\mathbf{h}_{\text{c}}\mathbf{h}_{\text{c}}^{H}$. In order to deal with the coupled optimization variables $\{\mathbf{x}^{\text{p}}\}$ and $\{\mathbf{W}\}$, we consider the penalty-based AO framework below.
III-A Radiation Pattern Design
With the given $\{\mathbf{x}^{\text{p}}\}$, the problem (P3) is reformulated as
$(\text{P}4)\ \max\limits_{\mathbf{W}}\quad$ $\text{Tr}(\mathbf{H}_{\text{s}}\mathbf{W})$ (11a)
@@ III-A Radiation Pattern Design
The problem (P4) is a SDP problem with rank-one constraints, which contain $2$ feasible linear constraints and $2$ matrix variables. According to [8, Eq. (32)], it readily knows the optimal general-rank solutions of (P4) satisfy
$\text{rank}(\mathbf{W})\leq\sqrt{2},$ (12)
which indicates that $\text{rank}(\mathbf{W})=1$. By applying the conclusion, the optimal radiation pattern can be obtained.
III-B Antenna Activation Position Optimization
To deal with the intractable expression of variable channels, we introduce auxiliary variables $\mathbf{U}$ and $\mathbf{V}$ to replace $\mathbf{H}_{\text{s}}$ and $\mathbf{H}_{\text{c}}$, respectively. Thus, we have the equality constraints $\mathbf{U}=\mathbf{H}_{\text{s}}$ and $\mathbf{V}=\mathbf{H}_{\text{c}}$. By relocating the equality constraint to the objective function and serving as a penalty term, the problem (P3) can be equivalently rewritten as
$(\text{P}5)\quad$ $\max\limits_{\mathbf{x}^{\text{p}},\mathbf{U},\mathbf{V}}\text{Tr}(\mathbf{U}\mathbf{W})-\frac{1}{2\varrho}\chi$ (13a)
s.t. $\text{Tr}(\mathbf{V}\mathbf{W})\geq\gamma_{\text{QoS}}\sigma^{2},$ (13b)
$\text{rank}(\mathbf{U})=1,\ \text{rank}(\mathbf{V})=1,$ (13c)
$\mathbf{U}\succeq\mathbf{0},\ \mathbf{V}\succeq\mathbf{0},$ (13d)
$\frac{1}{r_{\text{min,s}}^{2}}\leq\frac{\mathbf{U}_{[n,n]}}{\eta}\leq\frac{1}{r_{\text{max,s}}^{2}},$ (13e)
$\frac{1}{r_{\text{min,c}}^{2}}\leq\frac{\mathbf{V}_{[n,n]}}{\eta}\leq\frac{1}{r_{\text{max,c}}^{2}},$ (13f)
[…]
where $\chi=\|\mathbf{U}-\mathbf{H}_{\text{s}}\|_{F}+\|\mathbf{V}-\mathbf{H}_{\text{c}}\|_{F}$, and $\varrho$ denotes the scaling factor of the penalty terms. Here, $r_{\text{min,s}}$ and $r_{\text{max,s}}$ denote the minimum and maximum distances from an arbitrary pinching antenna to the sensing target. Also, $r_{\text{min,c}}$ and $r_{\text{max,c}}$ denote the minimum and maximum distances from an arbitrary pinching antenna to the communication user. To solve the problem (P5), we propose a penalty-based two-layer algorithm, which alternately optimizes $\{\mathbf{U},\mathbf{V}\}$ and $\{\mathbf{x}^{\text{p}}\}$ in the inner layer and update $\varrho$ in the outer layer.
III-B1 Inner layer iteration—subproblem with respect to $\{\mathbf{U},\mathbf{V}\}$
With the fixed $\{\mathbf{x}^{\text{p}}\}$, the problem (P5) is reduced to
$(\text{P}6)\ \max\limits_{\mathbf{U},\mathbf{V}}$ $\text{Tr}(\mathbf{U}\mathbf{W})-\frac{1}{2\varrho}\chi$ (14a)
[…]
To handle the rank-one constraint, we employ the difference-of-convex (DC) relaxation method [9] to rewrite the (13c) as
$\begin{cases}\Re(\text{Tr}(\mathbf{U}^{H}(\mathbf{I}-\mathbf{u}\mathbf{u}^{H})))\leq\varrho_{1},\\ \Re(\text{Tr}(\mathbf{V}^{H}(\mathbf{I}-\mathbf{v}\mathbf{v}^{H})))\leq\varrho_{2},\end{cases}\ i\in\{1,2\},$ (15)
where $\mathbf{u}$ and $\mathbf{v}$ represent the leading eigenvectors of $\mathbf{U}$ and $\mathbf{V}$. As a result, the problem (P6) can be transformed into
$(\text{P}7)\ \max\limits_{\mathbf{U},\mathbf{V},\varrho_{i}}\quad$ $\text{Tr}(\mathbf{U}\mathbf{W})-\frac{1}{2\varrho}\chi-\sum_{i=1}^{2}\varrho_{i}$ (16a)
s.t. $\varrho_{i}>0,\ i\in\{1,2\},$ (16b)
[…]
which is a convex problem and can be directly solved. Thus, the rank-one solution $\{\mathbf{U},\mathbf{V}\}$ can be obtained by iteratively solving the problem (P7).
III-B2 Inner layer iteration—subproblem with respect to $\{\mathbf{x}^{\text{p}}\}$
Note that the equality constraint $\mathbf{U}=\mathbf{H}_{\text{s}}$ and $\mathbf{V}=\mathbf{H}_{\text{c}}$ are equivalent to $\mathbf{u}=\mathbf{h}_{\text{s}}$ and $\mathbf{v}=\mathbf{h}_{\text{c}}$, where $\{\mathbf{u},\mathbf{v}\}$ denote the leading eigenvectors of $\{\mathbf{U},\mathbf{V}\}$. As a result, the problem (P6) can be transformed into
$(\text{P}8)\ \min\limits_{\mathbf{x}^{\text{p}}}\quad$ $\|\mathbf{u}-\mathbf{h}_{\text{s}}\|^{2}+\|\mathbf{v}-\mathbf{h}_{\text{c}}\|^{2}$ (17a)
[…]
It is easy to notice that $x_{n}^{\text{p}}$ and $x_{m}^{\text{p}}$ ($n\neq m$) are separated in the objective function but coupled in the constraint (8f), which motivates us to adopt the element-wise optimization framework. Therefore, with the fixed $\{x_{1}^{\text{p}},\cdots,x_{n-1}^{\text{p}},x_{n+1}^{\text{p}},\cdots,x_{N}^{\text{p}}\}$, the subproblem with respect to $x_{n}^{\text{p}}$ is given by
$(\text{P}9)\ \min\limits_{x_{n}^{\text{p}}}\quad$ $\varpi(x_{n}^{\text{p}})$ (18a)
s.t. $x_{n}^{\text{p}}\in[x_{n-1}^{\text{p}}+\Delta x,x_{n+1}^{\text{p}}-\Delta x],$ (18b)
where $\varpi(x_{n}^{\text{p}})$ is given by
$\varpi(x_{n}^{\text{p}})=\|\mathbf{u}_{[n]}-\mathbf{h}_{[n],\text{s}}\|^{2}+\|\mathbf{v}_{[n]}-\mathbf{h}_{[n],\text{c}}\|^{2}.$ (19)
@@ III-B2 Inner layer iteration—subproblem with respect to $\{\mathbf{x}^{\text{p}}\}$
III-B3 Outer layer iteration
@@
$\varrho=\varrho\bar{c},$ (20)
where $10<\bar{c}<$ 1 is the iteration coefficient of the penalty terms. The two-layer algorithm is summarized in Algorithm 1.
III-C Overall Algorithm
Algorithm 1 Two-layer algorithm.
1:  Initialize $\mathbf{x}^{\text{p}}$, $\mathbf{u}$, and $\mathbf{v}$. Set the convergence accuracy $\epsilon_{1}$, $\epsilon_{2}$, and $\epsilon_{3}$.
@@ III-C Overall Algorithm
4:         update $\{\mathbf{U},\mathbf{V}\}$ by iteratively solving the subproblem (P7) with an accuracy of $\epsilon_{1}$.
5:         update $\mathbf{x}^{\text{p}}$ by adopting element-wise exhaustive search.
@@
7:      $\varrho=\varrho\bar{c}$.
8:  until  $\varrho\leq\epsilon_{3}$.
Algorithm 2 Penalty-based AO algorithm.
1:  Parameter Initialization. Set a convergence accuracy $\epsilon_{4}$.
2:  repeat
3:      update $\{\mathbf{W}\}$ by solving the problem (P4).
4:      carry out Algorithm 1.
5:  until the objective value converges with an accuracy of $\epsilon_{4}$.
The proposed penalty-based AO algorithm is summarized in Algorithm 2, which is assured to converge at least to a stationary point solution. The computational complexity of Algorithm 2 mainly depends on solving the SDP problems (P4), (P7), and the one-dimensional exhaustive search. It is given by $\mathcal{O}\Big{(}\log(\frac{1}{\epsilon_{4}})\left(N^{3.5}+\log(\frac{1}{\epsilon_{3}})\log(\frac{1}{\epsilon_{2}})\big{[}\log(\frac{1}{\epsilon_{1}})N^{3.5}+N\bar{Q}\big{]}\right)\Big{)}$ [8], where $\bar{Q}$ represents the number of the quantization bits during the one-dimensional exhaustive search.

### v3

@@ II-C Problem Formulation
In this paper, we aim to maximize the illumination power $P(\theta_{\text{s}},r_{\text{s}})$ by designing the pinching beamformer, under the transmit power budget and communication QoS requirement, which is given by
$(\text{P}1)\quad$ $\max\limits_{\mathbf{x}^{\text{p}}}P_{\text{s}}$ (11a)
s.t. $R\geq R_{\text{QoS}},$ (11b)
$\mathbf{x}^{\text{p}}\in\mathcal{X},$ (11c)
where $R_{\text{QoS}}$ denotes the QoS requirement of the communication user. The problem (P1) is challenging to solve due to the quadratic objective function and the coupled variables.
@@ III Pinching Beamforming Optimization
In this section, we focus on the C&S transmission design by optimizing the pinching beamforming. To deal with the coupled optimization variables, a penalty-based AO algorithm is proposed, where $\{\mathbf{x}^{\text{p}}\}$ is optimized in an element-wise manner.
To facilitate the optimization, we can rewrite the problem (P1) as
$(\text{P}2)\ \max\limits_{\mathbf{x}^{\text{p}}}\quad$ $|\bm{\eta}^{H}\mathbf{v}(\mathbf{x}^{\text{p}})|^{2}$ (12a)
s.t. $|\bm{\eta}^{H}\mathbf{w}(\mathbf{x}^{\text{p}})|^{2}\geq\gamma_{\text{QoS}}\sigma^{2},$ (12b)
($,$ (12c)
where $\gamma_{\text{QoS}}=\frac{2^{R_{\text{QoS}}}-1}{P_{\text{T}}}$.
In order to deal with the intractable objective and constraints, we consider a penalty-based two-layer framework. To elaborate, we introduce auxiliary variables $\mathbf{\tilde{w}}$ and $\mathbf{\tilde{v}}$ to replace $\mathbf{w}(\mathbf{x}^{\text{p}})$ and $\mathbf{v}(\mathbf{x}^{\text{p}})$, respectively. Thus, we have the equality constraints $\mathbf{\tilde{w}}=\mathbf{w}(\mathbf{x}^{\text{p}})$ and $\mathbf{\tilde{v}}=\mathbf{v}(\mathbf{x}^{\text{p}})$. By relocating the equality constraint to the objective function and serving as a penalty term, the problem (P2) can be equivalently rewritten as
$(\text{P}3)\ \max\limits_{\mathbf{x}^{\text{p}},\mathbf{\tilde{w}},\mathbf{\tilde{v}}}\quad$ $|\bm{\eta}^{H}\mathbf{\tilde{v}}|^{2}-\frac{1}{2\varrho}\chi_{1}(\mathbf{x}^{\text{p}},\mathbf{\tilde{w}},\mathbf{\tilde{v}})$ (13a)
s.t. $|\bm{\eta}^{H}\mathbf{\tilde{w}}|^{2}\geq\gamma_{\text{QoS}}\sigma^{2},$ (13b)
$|\mathbf{\tilde{w}}_{[n]}|^{2}\leq\frac{1}{Nr_{\text{min,c}}^{2}},$ (13c)
$|\mathbf{\tilde{v}}_{[n]}|^{2}\leq\frac{1}{Nr_{\text{min,s}}^{2}},$ (13d)
($,$ (13e)
where $\chi_{1}(\mathbf{x}^{\text{p}},\mathbf{\tilde{w}},\mathbf{\tilde{v}})=\|\mathbf{\tilde{w}}-\mathbf{w}(\mathbf{x}^{\text{p}})\|+\|\mathbf{\tilde{v}}-\mathbf{v}(\mathbf{x}^{\text{p}})\|$ and $\varrho$ denotes the scaling factor of the penalty terms. Note that to avoid the infinite objective value, we introduce constraints (13c) and (13d), where $r_{\text{min,c}}=\sqrt{(r_{\text{c}}\sin\varphi_{\text{c}})^{2}+d^{2}}$ and $r_{\text{min,s}}=\sqrt{(r_{\text{s}}\sin\varphi_{\text{s}})^{2}+d^{2}}$ denote the lower bounds of the distances from an arbitrary pinching antenna to the communication user and target. The problem (P3) is equivalent to the problem (P1) as constraints (13c) and (13d) can be obtained from the (11c), which restricts pinching beamforming $\{\mathbf{\tilde{w}},\mathbf{\tilde{v}}\}$ to the feasible region.
To address the quadratic objective and constraints, we apply the SDR technique to rewrite the problem (P3) as follows.
$(\text{P}4)\quad$ $\max\limits_{\mathbf{x}^{\text{p}},\mathbf{\tilde{W}},\mathbf{\tilde{V}}}\text{Tr}(\bm{\eta}\bm{\eta}^{H}\mathbf{\tilde{V}})-\frac{1}{2\varrho}\chi_{2}(\mathbf{x}^{\text{p}},\mathbf{\tilde{W}},\mathbf{\tilde{V}})$ (14a)
s.t. $\mathbf{\tilde{W}}_{[n,n]}\leq\frac{1}{Nr_{\text{min,c}}^{2}},$ (14b)
$\mathbf{\tilde{V}}_{[n,n]}\leq\frac{1}{Nr_{\text{min,s}}^{2}},$ (14c)
$\text{Tr}(\bm{\eta}\bm{\eta}^{H}\mathbf{\tilde{W}})\geq\gamma_{\text{QoS}}\sigma^{2},$ (14d)
$\text{rank}(\mathbf{\tilde{W}})=1,\ \text{rank}(\mathbf{\tilde{V}})=1,$ (14e)
$\mathbf{\tilde{W}}\succeq\mathbf{0},\ \mathbf{\tilde{V}}\succeq\mathbf{0},$ (14f)
($,$ (14g)
where $\mathbf{W}(\mathbf{x}^{\text{p}})=\mathbf{w}(\mathbf{x}^{\text{p}})\mathbf{w}^{H}(\mathbf{x}^{\text{p}})$, $\mathbf{\tilde{W}}=\mathbf{\tilde{w}}\mathbf{\tilde{w}}^{H}$, $\mathbf{V}(\mathbf{x}^{\text{p}})=\mathbf{v}(\mathbf{x}^{\text{p}})\mathbf{v}^{H}(\mathbf{x}^{\text{p}})$, $\mathbf{\tilde{V}}=\mathbf{\tilde{v}}\mathbf{\tilde{v}}^{H}$, and $\chi_{2}(\mathbf{x}^{\text{p}},\mathbf{\tilde{W}},\mathbf{\tilde{V}})=\|\mathbf{\tilde{W}}-\mathbf{W}(\mathbf{x}^{\text{p}})\|_{F}+\|\mathbf{\tilde{V}}-\mathbf{V}(\mathbf{x}^{\text{p}})\|_{F}$. To solve the problem (P4), we propose a penalty-based AO algorithm, which alternately optimizes $\{\mathbf{\tilde{W}},\mathbf{\tilde{V}}\}$ and $\{\mathbf{x}^{\text{p}}\}$ in the inner layer and updates $\varrho$ in the outer layer.
III-1 Inner layer iteration—subproblem with respect to $\{\mathbf{\tilde{W}},\mathbf{\tilde{V}}\}$
With the fixed $\{\mathbf{x}^{\text{p}}\}$, the problem (P4) is reduced to
$(\text{P}5)\ \max\limits_{\mathbf{\tilde{W}},\mathbf{\tilde{V}}}$ $\text{Tr}(\bm{\eta}\bm{\eta}^{H}\mathbf{\tilde{V}})-\frac{1}{2\varrho}\chi_{2}(\mathbf{x}^{\text{p}},\mathbf{\tilde{W}},\mathbf{\tilde{V}})$ (15a)
@@ III-1 Inner layer iteration—subproblem with respect to $\{\mathbf{\tilde{W}},\mathbf{\tilde{V}}\}$
To handle the rank-one constraint, we introduce non-negative auxiliary variables $\{\varpi_{1},\varpi_{2}\}$ and employ the difference-of-convex (DC) relaxation method [9] to rewrite the (14c) as
$\begin{cases}\Re(\text{Tr}(\mathbf{\tilde{W}}^{H}(\mathbf{I}-\mathbf{\tilde{w}}_{\text{max}}\mathbf{\tilde{w}}_{\text{max}}^{H})))\leq\varpi_{1},\\ \Re(\text{Tr}(\mathbf{\tilde{V}}^{H}(\mathbf{I}-\mathbf{\tilde{v}}_{\text{max}}\mathbf{\tilde{v}}_{\text{max}}^{H})))\leq\varpi_{2},\end{cases}\ i\in\{1,2\},$ (16)
where $\mathbf{\tilde{w}}_{\text{max}}$ and $\mathbf{\tilde{v}}_{\text{max}}$ represent the eigenvectors corresponding to the maximum eigenvalues of $\mathbf{\tilde{W}}$ and $\mathbf{\tilde{V}}$, respectively. As a result, the problem (P5) can be transformed into
$(\text{P}6)\ \max\limits_{\mathbf{\tilde{W}},\mathbf{\tilde{V}},\varpi_{i}}\quad$ $\text{Tr}(\bm{\eta}\bm{\eta}^{H}\mathbf{\tilde{V}})-\frac{1}{2\varrho}\chi_{2}(\mathbf{x}^{\text{p}},\mathbf{\tilde{W}},\mathbf{\tilde{V}})-\sum_{i=1}^{2}\frac{1}{2\varrho_{i}}\varpi_{i}$ (17a)
s.t. $\varpi_{i}\geq 0,\ i\in\{1,2\},$ (17b)
@@
where $\varrho_{i}$ denotes the scaling factor of $\varpi_{i}$. The problem (P6) is a convex problem and can be directly solved. Thus, the rank-one solution $\{\mathbf{\tilde{W}},\mathbf{\tilde{V}}\}$ can be obtained by carrying out the Algorithm 1.
III-2 Inner layer iteration—subproblem with respect to $\{\mathbf{x}^{\text{p}}\}$
Note that the equality constraint $\mathbf{\tilde{W}}=\mathbf{W}(\mathbf{x}^{\text{p}})$ and $\mathbf{\tilde{V}}=\mathbf{V}(\mathbf{x}^{\text{p}})$ are equivalent to $\mathbf{\tilde{w}}=\mathbf{w}(\mathbf{x}^{\text{p}})$ and $\mathbf{\tilde{v}}=\mathbf{v}(\mathbf{x}^{\text{p}})$. As a result, the problem (P6) can be transformed into
$(\text{P}7)\ \min\limits_{\mathbf{x}^{\text{p}}}\quad$ $\|\mathbf{\tilde{w}}-\mathbf{w}(\mathbf{x}^{\text{p}})\|+\|\mathbf{\tilde{v}}-\mathbf{v}(\mathbf{x}^{\text{p}})\|$ (18a)
@@ III-2 Inner layer iteration—subproblem with respect to $\{\mathbf{x}^{\text{p}}\}$
It is easy to notice that $x_{n}^{\text{p}}$ and $x_{m}^{\text{p}}$ ($n\neq m$) are separated in the objective function but coupled in the constraint (11c), which motivates us to adopt the element-wise optimization framework. Therefore, with the fixed $\{x_{1}^{\text{p}},\cdots,x_{n-1}^{\text{p}},x_{n+1}^{\text{p}},\cdots,x_{N}^{\text{p}}\}$, the subproblem with respect to $x_{n}^{\text{p}}$ is given by
$(\text{P}8)\ \min\limits_{x_{n}^{\text{p}}}\quad$ $\left|\mathbf{\tilde{w}}_{[n]}-\frac{e^{-\jmath(\frac{2\pi}{\lambda}r_{n}^{\text{c}}(r_{\text{c}},\varphi_{\text{c}})+\theta_{n})}}{\sqrt{N}r_{n}^{\text{c}}(r_{\text{c}},\varphi_{\text{c}})}\right|$
$+\left|\mathbf{\tilde{v}}_{[n]}-\frac{e^{-\jmath(\frac{2\pi}{\lambda}r_{n}^{\text{s}}(r_{\text{s}},\varphi_{\text{s}})+\theta_{n})}}{\sqrt{N}r_{n}^{\text{s}}(r_{\text{s}},\varphi_{\text{s}})}\right|$ (19a)
s.t. $x_{n-1}^{\text{p}}+\Delta x\leq x_{n}^{\text{p}}\leq x_{n+1}^{\text{p}}-\Delta x,$ (19b)
$\frac{-L}{2}\leq x_{n}^{\text{p}}\leq\frac{L}{2},$ (19c)
@@
III-3 Outer layer iteration
@@
$\varrho=\varrho\bar{c}_{2},$ (20)
where $10<\bar{c}_{2}<$ 1 is the iteration coefficient of the penalty terms. The penalty-based AO algorithm is summarized in Algorithm 2.
Algorithm 1 Iterative algorithm for rank-one solution.
1:  Initialize $\mathbf{\tilde{v}}_{\text{max}}$, and $\mathbf{\tilde{w}}_{\text{max}}$. Set a convergence accuracy $\epsilon_{1}$.
2:  repeat
3:      update $\{\mathbf{\tilde{W}},\mathbf{\tilde{V}},\varpi_{i}\}$ by solving the problem (P6).
4:      update the eigenvectors $\{\mathbf{\tilde{w}}_{\text{max}},\mathbf{\tilde{v}}_{\text{max}}\}$.
5:      update $\varrho_{i}=\varrho_{i}\bar{c}_{1}$ ($10<\bar{c}_{1}<$ 1).
6:  until  $\sum_{i=1}^{2}\varpi_{i}$ falls below a threshold of $\epsilon_{1}$.
Algorithm 2 Penalty-based AO algorithm.
1:  Parameter Initialization. Set the convergence accuracy $\epsilon_{2}$ and $\epsilon_{3}$.
@@
4:         update $\{\mathbf{\tilde{w}},\mathbf{\tilde{v}}\}$ by carrying out Algorithm 1.
5:         update $\mathbf{x}_{\text{p}}$ via the element-wise optimization.
@@
7:     update $\varrho=\varrho\bar{c}_{2}$ ($10<\bar{c}_{2}<$ 1).
8:  until $\|\mathbf{\tilde{W}}-\mathbf{W}(\mathbf{x}^{\text{p}})\|_{F}+\|\mathbf{\tilde{V}}-\mathbf{V}(\mathbf{x}^{\text{p}})\|_{F}\leq\epsilon_{3}$.
The proposed penalty-based AO algorithm is summarized in Algorithm 2, which is assured to converge at least to a stationary point solution. The computational complexity of Algorithm 2 mainly depends on solving the SDP problems (P6) and the one-dimensional exhaustive search. It is given by $\mathcal{O}\Big{(}\log(\frac{1}{\epsilon_{3}})\log(\frac{1}{\epsilon_{2}})\big{[}\log(\frac{1}{\epsilon_{1}})N^{3.5}+N\bar{Q}\big{]}\Big{)}$ [10], where $\bar{Q}$ represents the number of the quantization bits during the one-dimensional exhaustive search.